OpenStudy (anonymous):

how do you simplify ln e^4 ?

OpenStudy (anonymous):

C'mon now. You've seen this at least 4 times already... \[ln(e^4) = k \iff e^k = e^4\] So what is k?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Well thanks Nancy, but I wasn't really asking you ;p

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

sorry it took so long, but i forgot